# Quantum field theory on curved manifolds

### v1

@@ Abstract
Given a manifold of a system with internal degrees of freedom, such as Lorentz symmetry or gauge symmetry, the “curvature” is defined for the manifold. If one defines the local vacuum in the tangent space of the manifold, one can define a local mapping in the vicinity of the contact point, which is nothing but the Bogoliubov transformation. The curvature of the electromagnetic field gives the Schwinger effect, and the curvature of the base space introduces the local Unruh effect, which realizes Hawking radiation if applied on the Black hole horizon. To show how the Bogoliubov transformation appears on the manifold and why the local calculation is crucial there, we consider the Schwinger effect with a slowly varying electric field. Then, we show what happens in the Unruh effect if the acceleration is not a constant parameter.
@@ 1 Introduction
When the curvature of a manifold is given by a non-zero constant, a dynamical element of particle generation may be added to the system, even though the manifold itself seems to be static. The Schwinger effect[1] and Hawking radiation[2] are typical examples of such phenomena. Particle production in spacetime with curvature has been treated by various methods[3]. In this paper, we focus on the most primitive method for calculating Bogoliubov transformations: analysis with field equations. In this case, the Bogoliubov transformations are realized as the Stokes phenomena of the differential equation. In the study of particle production, when it is analyzed using field equations, it is customary to introduce asymptotic states (flat space at infinity in time or space) to define the “vacuum” and the particle number there. One might argue that the loss of locality is inevitable in such calculation, but at the same time, it is hard to believe that loss of locality is an essential feature of the phenomena. Indeed, if particle production is to be a heat bath, the loss of Markov property seems to be fatal for the analysis. In principle, it is also strange that the calculations cannot be completed at the place where the particles are created, without using the distant past or the distant future. Therefore, the loss of locality is clearly due to problems with the computational procedures. Turning to mathematics, gauge theory and general relativity are constructed by differential geometry and manifolds based on local analysis[4, 5, 6]. It is therefore natural to ask whether, using mathematical concepts based on local analysis, these problems can be solved without assuming asymptotic states at a distance, even if they are analyzed using field equations. In particular, in mathematical concepts of manifolds, tangent spaces appear naturally. These tangent spaces have properties quite similar to the asymptotic states. The reasons why a tangent space can be used as a vacuum are explained in detail in this paper. Since the tangent space is an essential feature of manifolds and differential geometry, and it has the properties required to define the vacuum in it, then it is very natural to assume that the vacuum has to be defined in the tangent space rather than for the asymptotic states. The aim of this paper is to rethink particle production on manifolds based on mathematical concepts. The best-known tangent space in physics is the local inertial system of general relativity. In the electromagnetic theory, it might be strange to define a tangent space, but indeed the tangent space of the principle bundle is commonly used in mathematics to define the curvature of the electromagnetic field. Our perspective is that particle production on manifolds, such as the Schwinger and the Unruh effects should be discussed in a unified way by using the same basic concepts of the differential geometry and manifolds. The crucial difference between the Unruh effect[7] and the Schwinger effect will be made clear in this paper.
@@
When attempting to solve locally what has long been solved globally in physics, it is necessary to reconsider various mathematical properties that have previously been ignored. One might wonder why such mathematical concepts are required even in situations where solutions by path integrals are easily given. On this point, it would be important to emphasize here that path integral and Feynman diagrams are so cleverly constructed that a user can sometimes get the right result without having to worry about such mathematical concepts.
@@
Section 2 of this paper describes the basic concepts of differential geometry and of manifolds, and the introduction of the symbols that are used in this paper. Section 3 describes the Schwinger effect when the electric field is constant or time-dependent. For fermions, it looks like the Landau-Zener transition[8]. By considering the case where the electric field is weakly depending on time, we show why the vacuum must be defined in the tangent space. This is the simplest example that clearly demonstrates that finding a naive global exact solution to an equation cannot always yield a physical solution. Section 4 discusses a local analysis of the Unruh effect and Hawking radiation. Again, as in the case of the Schwinger effect, the importance of defining the vacuum in the tangent space and the importance of local analysis is confirmed by considering the case of gradually varying acceleration.
@@ 2 An introduction to differential geometry and manifolds for particle production
First, we briefly describe the basic concepts of the manifold. A manifold is a fundamental concept in mathematics. It is defined as a topological space that locally resembles Euclidean space. This means that for every point in the manifold, there exists a neighborhood that can be mapped homeomorphically (i.e., through a continuous, bijective function with a continuous inverse) to an open subset of $\mathbb{R}^{n}$, where $n$ is the dimension of the manifold. For physics, this means that in principle the manifold will always have the required structure that is needed to define a local vacuum. [^2] To understand this more clearly for the Schwinger effect, we need to rethink tangent spaces after introducing gauge symmetries and matter fields. This local structure was designed from purely mathematical considerations. We will list here the most important definitions (and notations) of the manifolds that are convenient for our later discussions.
@@
    • A chart is a pair $(U_{i},\varphi_{i})$.
@@
Then, a tangent bundle is defined by
@@
Also, the cotangent space at the point $p$, denoted as $T^{*}_{p}M$, is defined as the dual space of the tangent space $T_{p}M$. In this paper, both the tangent space and the cotangent space can simply be referred to as the tangent space if there is no possibility of confusion. For the sake of the explanation that follows, we are going to start with an intuitive explanation about defining the vacuum in the tangent space. We are now going to construct a field theory on $M$ as manifolds: the field theory constructed on $M$ is, of course, also defined at $p\in M$. Then, the theory defined on $p$ is naturally extended into the tangent space. A key feature of the theory extended to tangent spaces is that curvatures vanish. This is trivial with respect to the spacetime curvature, as the tangent space is obviously flat, but it should not be trivial with respect to gauge field curvature, as the gauge field curvature may not vanish in the flat spacetime. The vanishing curvature of the gauge field is understood by the local trivialisation that will be introduced in the definition of the fibre bundle in the following. Although a tangent space can be defined on any space defined as a manifold, here the vacuum is defined in the tangent space of $M$, as explained above.
Above, we have described tangent vector bundles to introduce the concept of bundles. It seems obvious that one can consider similar vector bundles in general (not necessarily for tangent vectors of $M$). This is the idea behind the fibre bundle. The coordinate functions defined in tangent vector bundles naturally had a direct product structure. [^3] Intuitively, the local direct product structure gives a point where one can start the discussion with a global symmetry (because they are given by the direct product) when describing a local symmetry. When considering general fibre bundles, the direct product structure has to be introduced by hand. To clarify the notation used, we refer to the definitions below.
@@
Thanks to the introduction of a local direct product structure in the fibre, connections can be defined in a natural way. At this point, the introduction of a direct product structure may not seem to make sense as physics, but when one considers the role that local inertial systems actually play in general relativity, one has to consider that it does make sense as physics. Since in mathematics general relativity and gauge theories are described in a unified way, it seems to be natural to treat the two in a unified way. We believe that essential differences are highlighted only when analyses are held on the same foundations as far as possible.
Before introducing the connection, we should mention a scalar field used in the field theory of physics, which appears as a section of $E$ of a manifold. A “section” of a manifold will need some explanation. If we think of the elementary function $y=f(x)$ as a map $\mathbb{R}\rightarrow\mathbb{R}$, this function cuts $\mathbb{R}$ at the destination and gives a single value; if we draw a graph of $y=f(x)$ on the $xy$ space, its appearance will look like considering a section of the $xy$ space. Literally, a scalar field of the field theory corresponds to a section of the corresponding mapping. More abstractly, suppose that we are given some kind of mapping; If we give this mapping a concrete form, we are looking at a section. For cases where the mapping has internal degrees of freedom, a section is seen as one concrete form is selected. In physics, selecting the frame of a particular observer in special relativity is a section of the frame bundle.
@@
Above, fibres have been introduced locally by means of local trivialization, and fibres on $U_{i}$ are laminated together to form a fibre bundle. When local trivialization varied from place to place, it was necessary to connect them by using the transformation. Therefore, it is very natural to ask whether the fibre bundle that is obtained in this simple way really has the correct differential structure. If one actually prepares a vector bundle [^4] The vector bundle has any dimension and is not always supposed to be the tangent vector bundle. and its section (e.g, a scalar field) and simply takes the derivative, it can be seen that the simple derivative is not covariant. Then, connections are introduced to solve this problem. This is called a covariant derivative. This connection is necessary for the differential geometry because fibre bundles are locally trivialized and then laminated together using transformations. If trivialization is possible in global, then inevitably the connections will disappear. [^5] On the other hand, if the section is defined for a moving frame, the connections do not vanish for the observer, even if the fibres are globally trivial on the manifold. The same is true for a “moving gauge” of the Schwinger effect. This point will be explained in more detail later. For the same reason, the value of the connection is supposed to vanish at the contact point of the tangent space.
@@
which translates the same way as the original section (i.e, the scalar field $\phi$) of the vector bundle. A simple explanation is that since the two sections $\phi$ and $\nabla\phi$ transform in the same way, they can be laminated in the same way. This is the requirement for the consistency of the differential geometry. Here $A$ is called a connection or a gauge field. Although the transformation of the scalar field is trivial under the coordinate transformation, the effect of gravity is incorporated in a natural way as $d\phi$ is also a section of the frame bundle. The frame bundle is described below.

### v5

@@ Abstract
This paper discusses how particle production from the vacuum can be explained by local analysis when the field theory is defined by differential geometry on curved manifolds. We have performed the local analysis in a mathematically rigorous way, respecting the Markov property. The exact WKB is used as a tool for extracting non-perturbative effect from the local system. After a serious application of the differential geometry and the exact WKB to particle production, we show that entanglement does not appear in the Unruh effect as far as the standard formulation by the differential geometry is valid. This result should not be attributed to a consistency problem between the “entanglement state” and the “standard field theory by differential geometry”, but to the fact that the conventional calculation of the Unruh effect is done by extrapolation which is not consistent with the differential geometry. The situation is similar to that of the Dirac monopole, but topology is not relevant and the basis for building field theories in differential geometry is strongly involved.
@@ 1 Introduction
The knowledge gained by constructing field theory by differential geometry is significant. The most famous is what happned in 70’s to the solution of the Dirac monopole[1, 2], where the differential geometry revealed that at least two solutions are involved for the Dirac monopole. The solution is famous in topology, but the mathematician’s deep understanding of local analysis in differential geometry is at the root of this story. Similar discoveries are likely to continue in the future, as differential geometry can involve various mathematical concepts to construct the field theory in a variety of sophisticated ways. In this paper, the Schwinger effect[3] and the Unruh effect[5, 4, 6] are analysed using the standard formulation of the field theory in terms of differential geometry. We paid particular attention to the Markov property and taken great care to ensure that the computation is localised and consistent with differential geometry.
@@
When the curvature of a manifold is given by a non-zero constant, a dynamical element of particle generation may be added to the system, even though the manifold itself seems to be static. The Schwinger effect[3] and Hawking radiation[4] are typical examples of such phenomena. Particle production in spacetime with curvature has been treated by various methods[6], but we are dissatisfied with conventional calculations because of loss of locality. In this paper, we focus on the most primitive method for calculating Bogoliubov transformations: analysis with field equations. In this case, the Bogoliubov transformations are realized as the Stokes phenomena of the differential equation. The main reason for us to stick to differential equations is that they are compatible with differential geometry. In the study of particle production, when it is analyzed using field equations, it is customary to introduce asymptotic states (flat space at infinity in time or space) to define the “vacuum” and the particle number there. One might argue that the loss of locality is inevitable in such calculation, but at the same time, it is hard to believe that loss of locality is an essential feature of the phenomena. Indeed, if particle production is to be a heat bath, the loss of Markov property seems to be fatal for the analysis. In principle, it is also strange that the calculations cannot be completed at the place where the particles are created, without using the distant past or the distant future. Therefore, the loss of locality is clearly due to “for convenience”, which is the problem with the computational procedures of the physics side. Turning to mathematics, gauge theory and general relativity are constructed by differential geometry and manifolds, where local consistency is quite rigorous[7, 8, 9]. It is therefore natural to ask whether, using mathematical concepts, these problems can be solved without assuming asymptotic states at a distance, or without introducing extrapolation. It could be difficult to notice the unnaturalness we are discussing about when one is accustomed to ordinary calculations in physics. However, we strongly insist that loss of locality is a serious problem that should be resolved by mathematics. In particular, in mathematical concepts of manifolds, tangent spaces appear naturally. These tangent spaces have properties quite similar to the asymptotic states but appear in local. To explain in more detail, an operation called local trivialization is required at the point of contact with the tangent space, and this operation ensures that the connection is always zero at that point. The most familiar tangent space in physics will be the local inertial system, which is of course important in physics.222 To be more precise, the “inertial system” mentioned above is the tangent space with the observer’s (subjective) inertial frame. (The “frame” in the differential geometry is defined on the tangent space.) In physics, it is known that the subjective inertial frame causes Thomas precession[12]. The reasons why a tangent space can be used as a vacuum are explained in detail in this paper, considering a simple example where the vacuum must be defined in the local tangent space. Indeed, considering how to introduce Lorentz symmetry into the local system, it is obvious that a vacuum must be defined in the tangent space even if it is defined as an asymptotic state at a distance. This point will be clarified in the discussion of charts and frames. Since both Lorentz and gauge symmetries are treated equivalently in the framework of differential geometry, it is quite natural to assume that the tangent space to the principal bundle of a gauge transformation has the same physical meaning.
Note that the particle production discussed here is not an “overtly dynamical phenomenon” like prereheating after inflation, but rather a “phenomenon that causes dynamical particle production despite its apparent static state”. We will explain later that for the latter case a non-trivial treatment regarding gauge and Lorentz symmetries is required, which is difficult to be described without the help of differential geometry. In this paper, we will try to explain mathematical ideas as intuitive as possible, but since the concepts of the mathematics are rather complex, at least a basic knowledge of a typical differential geometry course is required to understanding the calculations correctly. Unlike differential geometry, which could be familiar as an undergraduate mathematics course, it would be very difficult to understand the exact WKB without reading any of the references. The serious difficulty with the exact WKB is that it looks so much like a regular WKB approximation that it is sometimes very hard to imagine how different the mathematics behind it is.
The aim of this paper is to rethink locality of particle production, keeping the discussions as faithful as possible to mathematics. The main obstacle to the local description of non-perturbative particle production has been to find a local analysis of non-perturbative effect, which is now solved by the exact WKB in mathematics. The best-known tangent space in physics would be the local inertial system of general relativity. In the electromagnetic theory, it might be strange to define a tangent space, but indeed the tangent space of the principle bundle is commonly used in mathematics to define the curvature of the electromagnetic field. Our perspective is that particle production on manifolds, such as the Schwinger and the Unruh effects should be discussed in a unified way by using the same basic concepts of the differential geometry and manifolds. The crucial difference between the Unruh effect[5] and the Schwinger effect will be made clear in this paper, by using local calculations faithful to mathematics.
@@
When attempting to solve locally what has long been solved globally in physics, it is necessary to reconsider various mathematical properties that have previously been ignored. One might wonder why such mathematical concepts are required even in situations where solutions by path integrals are easily given. On this point, it would be important to emphasize here that path integral and Feynman diagrams are so cleverly constructed that a user can sometimes get the right result without having to worry about such mathematical concepts. If the computation by path integrals is locally complete, our analysis will not lead to any new conclusions.
Section 2 of this paper describes the basic concepts of differential geometry and of manifolds, and the introduction of the symbols that are used in this paper. Since differential geometry has almost never been seriously discussed in the analysis of the Unruh effect, we will be particularly careful to explain which aspects we will focus on. As the concept of manifolds is not explained from scratch, the reader who is not familiar with the mathematical approach is advised to read the references.
@@
Section 3 describes the Schwinger effect when the electric field is constant or time-dependent. For fermions, it looks like the Landau-Zener transition[11]. By considering the case where the electric field is weakly depending on time, we show why the vacuum must be defined in the local tangent space. This is the simplest example that clearly demonstrates that finding a naive global exact solution to an equation cannot always be a correct answer. Note that the basic concept explaining that a single solution cannot cover the whole space is the same as that of the Dirac monopole, although it has nothing to do with topology. In order to understand the causes of this non-trivial situation, it is necessary to take a closer look at the treatment of gauge and Lorentz symmetries in differential geometry, especially at the necessity of local trivialization. At the time of the Dirac monopole, much attention was paid to topology, but in the root of the discussions there has been the basic procedures in finding local solutions in differential geometry.
Section 4 discusses a local analysis of the Unruh effect and Hawking radiation. Unlike the Unruh effect, where entangled pair production is expected at distant wedges, Hawking radiation is a pair production localised at the horizon. Therefore, we can calculate local particle production of Hawking radiation rigorously using differential geometry and the exact WKB, without indication of any discrepancy between the standard result.
@@ 2 An introduction to differential geometry and manifolds for particle production
First, we briefly describe the basic concepts of differential geometry and manifolds for the field theory. For our argument in this paper, a clear distinction between “charts” and “frames” is most important. Therefore, the main purpose of this introduction is to explain clearly why they are distinguished in mathematics and what are the implications for physics.
We know that in physics the vacuum is exactly the same for observers in any frame. In mathematics, the Lie algebra and bundles are used to define the situation. However, the Lie algebra and the bundles are not defined for “charts”. As the Rindler coordinate system is a “chart”, it was (at least in principle) unnatural to define a “vacuum” in the Rindler chart to discuss the Bogoliubov transformation. Furthermore, the “moving frame” of an observer has its applicable range. The range is estimated as $\propto 1/a$ in Ref.[12], where $a$ is the rate of acceleration. The extrapolation beyond this range is not rigorous, but such extrapolation is normally used in the conventional calculation of the Unruh effect. Therefore, there is a need for a method that allows rigorous calculations to be made in local. Our idea is to introduce the exact WKB for the local analysis, using the vierbeins. Without understanding the unnaturalness described here, it will be hard to understand the following arguments.
Although we will try to provide an overview as intuitive as possible, a basic knowledge of ordinary differential geometry is essential for a proper understanding of the subject. The standard description used here can be found in Ref.[7, 8, 9], but in principle different (and more complicated) extensions of manifolds could be possible in the description of the field theory[10]. The mathematical concepts associated with field theory are not always confined to the basic scenario presented here.
A manifold is a fundamental concept in mathematics. It is defined as a topological space that locally resembles Euclidean space. This means that for every point in the manifold, there exists a neighborhood that can be mapped homeomorphically (i.e., through a continuous, bijective function with a continuous inverse) to an open subset of $\mathbb{R}^{n}$, where $n$ is the dimension of the manifold. For physics, this implies that in principle the manifold will always have the required structure that is needed to define a local vacuum. [^3] To understand this more clearly for the Schwinger effect, we need to rethink tangent spaces after introducing gauge symmetries and matter fields. The local structure of the differential geometry was designed from purely mathematical considerations. We will list here the most important definitions (and notations) of the manifolds that are convenient for our later discussions. In field theory, it is not possible to explain everything with just one manifold. Such complex structure is important for introducing gauge and Lorentz symmetries.
@@
    • A chart is a pair $(U_{i},\varphi_{i})$. (See fig.1.)
@@
Figure 1: A local coordinate function $\varphi_{i}$ is shown as a homeomorphism from $U_{i}\in M$ to $V_{i}\in\mathbb{R}^{n}$. A chart is a pair $(U_{i},\varphi_{1})$. The Rindler coordinate is a typical chart on flat spacetime. The “chart” must be discriminated from the “frame”.
@@
Then, a tangent bundle is defined by [^4] The tangent bundle is also a $2n$-dimensional manifold.
@@
Suppose that $\varphi_{i}(p)$ is the coordinate function $\{x^{\mu}(p)\}$ of $U_{i}$. In the “coordinate basis”, $T_{p}M$ is spanned by $\{e_{\mu}\}=\{\partial/\partial x^{\mu}\}$, while the “non-coordinate bases” is explained as
where the coefficients $e_{\alpha}^{\mu}$ are called vierbeins. Note that the Lie brackets can be introduced only for the non-coordinate bases where the vierbeins play an essential role, as is shown in Fig.2. The figure shows that the vacuum must be defined in the tangent space, even if it is defined as an asymptotic state.
Figure 2: The tangent space $T_{p}M$ is spanned by $\{e_{\mu}\}=\{\partial/\partial x^{\mu}\}$, where $\{x^{\mu}\}$ is the coordinate function of $U_{i}$. As is shown in Eq.(2.5) and the above picture, the vierbeins are essential for defining the Lie algebra. This means that the vacuum must be defined in the tangent space as far as the vacuum respects the Lorentz symmetry, even if it is placed at far distance as to describe the asymptotic state. The frame of an accelerating observer is called a moving frame, as it (the vierbeins) is changed by the Lorentz transformation time to time.
Also, the cotangent space at the point $p$, denoted as $T^{*}_{p}M$, is defined as the dual space of the tangent space $T_{p}M$. In this paper, both the tangent space and the cotangent space can simply be referred to as the tangent space if there is no possibility of confusion. For the sake of the explanation that follows, we are going to start with an intuitive explanation about defining the vacuum in the tangent space. We are now going to construct a field theory on $M$ as manifolds [^5] We need many kinds of manifolds at the same time to explain the basic properties of the field theory.: the field theory constructed on $M$ is, of course, also defined at $p\in M$. Then, the theory defined on $p$ is naturally extended into the tangent space. A key feature of the theory extended in the tangent space is that curvatures vanish. This is trivial with respect to the spacetime curvature, as the tangent space is obviously flat, but it should not be trivial with respect to gauge field curvature, as the gauge field curvature may not vanish in the flat spacetime. The vanishing curvature of the gauge field is understood if the gauge theory in the tangent space (of the spacetime) is also described by the tangent space of the principal bundle of the gauge. Thus, by treating both Lorentz and gauge symmetries equally, the local vacuum can be defined as having no curvature in either sense. Since a tangent space can be defined on any space defined as a manifold, here the vacuum is defined using these tangent spaces. What will be important in our calculations is that the connection is always zero at the contact point with the tangent space, where the local vacuum is defined.
Above, we have described tangent vector bundles to introduce the concept of bundles. It seems obvious that one can consider similar vector bundles in general (not necessarily for tangent vectors of $M$). This is the idea behind the fibre bundle. The coordinate functions defined in tangent vector bundles naturally had a direct product structure. [^6] Intuitively, the local direct product structure gives a point where one can start the discussion with a global symmetry (because they are given by the direct product) when describing a local symmetry. This procedure is one that we often see in field theory, whether or not there is a strict definition of local trivialization. When considering general fibre bundles, the direct product structure has to be introduced by hand. To clarify the notation used, we refer to the definitions below.
@@
Thanks to the introduction of a local direct product structure in the fibre, connections can be defined in a natural way. At this point, the introduction of a direct product structure may not seem to make sense as physics. However, when one considers the role that local inertial systems actually play in general relativity, one has to consider that it does make sense as physics. Since in mathematics, general relativity and gauge theories are described in a unified way, it seems to be natural to treat the two in a unified way. Here, the “gauge field with curvature” yields the Schwinger effect, while the “spacetime with curvature” yields the Unruh effect and Hawking radiation. We believe that essential differences are highlighted only when analyses are held on the same foundations as far as possible.
Before introducing the connection, we should mention a scalar field used in the field theory of physics, which appears as a section of $E$ of a manifold. A “section” of a manifold will need some explanation. If we think of the elementary function $y=f(x)$ as a map $\mathbb{R}\rightarrow\mathbb{R}$, this function cuts $\mathbb{R}$ at the destination and gives a single value; if we draw a graph of $y=f(x)$ on the $xy$ space, its appearance will look like considering a section of the $xy$ space. Literally, a scalar field of the field theory corresponds to a section of the corresponding mapping. More abstractly, suppose that we are given some kind of mapping; If we give this mapping a concrete form, we are looking at a section. For cases where the mapping has internal degrees of freedom, a section is seen as one concrete form is selected. In physics, selecting the frame of a particular observer in special relativity gives a section of the frame bundle. For an accelerating observer, the observer’s frame moves on the frame bundle. This non-trivial section is called the moving frame. The range of applicability of the moving frame has been calculated in Ref.[12]. It is only consistent in the range $\propto a^{-1}$, where $a$ is the accelerating rate of the observer. This point is very important, since the conventional calculation extrapolate well beyond this range and this may have led to an entanglement that should not have occurred. Hypotheses of this kind can only be tested by means of local calculations without extrapolation.
@@
Above, fibres have been introduced locally by means of local trivialization, and the fibres are laminated together to form a fibre bundle. When local trivialization varied from place to place, it was necessary to connect them by using the transformation. Therefore, it is very natural to ask whether the fibre bundle that is obtained in this simple way really has the correct differential structure. If one actually prepares a vector bundle [^7] The vector bundle has any dimension and is not always supposed to be the tangent vector bundle. and its section (e.g, a scalar field) and simply takes the derivative, it can be seen that the simple derivative is not covariant. Then, connections are introduced to solve this problem. This is called a covariant derivative. This connection is necessary for the differential geometry because fibre bundles are locally trivialized and then laminated together using transformations. If trivialization is possible in global, then inevitably the connections will disappear. [^8] On the other hand, if the section is defined for a moving frame, the connections do not vanish for the observer, even if the fibres are globally trivial on the manifold. The same is true for a “moving gauge” of the Schwinger effect. This point will be explained in more detail later. For the same reason, the value of the connection is supposed to vanish at the contact point of the tangent space. The reason why non-trivial connections appear in the equations of an observer in accelerated motion in flat space is very easy to understand from a differential geometry point of view. This is because, although the space-time itself is flat, the observer’s equation sees a non-trivial section of the bundle.
@@
which translates the same way as the original section (i.e, the scalar field $\phi$) of the vector bundle. A simple explanation is that since the two sections $\phi$ and $\nabla\phi$ transform in the same way, they can be laminated in the same way. This is the requirement for the consistency of the differential geometry. Here $A$ is called a connection or a gauge field. Although the transformation of the scalar field is trivial under the coordinate transformation, the effect of gravity is incorporated in a natural way as $d\phi$ is also a section of the frame bundle. The frame bundle will be described later.